How many Images can form in double mirrors
1
Expert's answer
2019-12-02T09:52:33-0500

The formula to find the number of images n formed between two plane mirrors inclined at angle "\theta" between them is as follows:


"n=\frac{360}{\theta}-1"

"For\ \theta\to 0,\ n\to \infty"

Answer: infinity.